OpenStudy (anonymous):

Write the sum or difference in simplest form. Plasma makes up 11/20 of your blood, and blood cells make up the other 9/20. How much more of your blood is plasma than blood cells? A 1/20 B 1/10 C 1 D 11/20

OpenStudy (anonymous):

11/20-9/20=2/20=1/10 b